This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for beam experts, Page 1 of 3  NA61/SHINE logo
ID Date Author Subject Text Attachments
  297   Wed, 5.08.2026 15:04:13 K. WojcikPb test run settings    
  296   Wed, 5.08.2026 15:03:03 K. WojcikFinal settings for Be runBeam counters:

[LIST]
  
  295   Wed, 29.07.2026 19:40:38 M. QuintieriBeam settings for Be run   printscreen_295_20260729-194057.pngprintscreen_295_20260729-195658.pngprintscreen_295_20260729-200026.png 
  294   Thu, 23.07.2026 20:20:45 Aleksandr DmitrievSettings for p + LNT Long @ 15GeV/c   printscreen_294_20260723-201951.png 
  293   Thu, 16.07.2026 14:04:09 MQA new trim scan after beam tuning by Govanni   printscreen_293_20260716-140312.pngprintscreen_293_20260716-140450.pngprintscreen_293_20260716-143537.pngprintscreen_293_20260719-140021.png 
  292   Tue, 14.07.2026 23:10:04 Yuliia Balkova, Bartosz MaksiakCount recoveryYB: Around 19:20, Giovanni came to try to
focus our beam better. After half an hour
of unsuccessful attempts, he resigned and
 Screenshot_from_2026-07-14_23-09-09.pngScreenshot_from_2026-07-14_22-58-15.pngScreenshot_from_2026-07-14_20-26-15.pngprintscreen_292_20260715-175745.png 
  291   Sun, 12.07.2026 16:42:56 BM, AD, YBTrim scan 15 GeV/c   printscreen_290_20260712-164203.pngprintscreen_291_20260712-164930.png 
  290   Sun, 12.07.2026 14:32:19 Michael Quintierip@15 GeV CEDAR ScanCentral value of proton peak: 8.85 bar.  printscreen_290_20260712-143325.pngprintscreen_290_20260712-143351.pngprintscreen_290_20260712-143620.png 
  289   Fri, 10.07.2026 11:20:17 S. PulawskiS3 at p13GeV/c preparation scans   printscreen_289_20260710-112157.pngprintscreen_289_20260710-113645.png 
  288   Thu, 19.03.2026 14:25:49 K. Wojcik, A. Shuklapp run 2026 - after beam retuning:Beam was retuned upstream, so counters needed
to be aligned. It improved T2/T1. CEDAR was
also realigned. No new pressure scan needed.
 S5_scan_X_19-03-2026.pngS5_scan_Y_19-03-2026.pngCESAR_GUI_10.1.26_19-Mar-26_13-46-13.pngCESAR_GUI_10.1.26_19-Mar-26_13-19-25.pngCESAR_GUI_10.1.26_19-Mar-26_15-12-11.png 
  287   Sat, 14.03.2026 15:45:43 K. Wojcikpp run 2026: starting physics, field ON setupCEDAR:
  - pressure: 10.371 [bar]
  - diaphragm: 0.149 [mm]
 S5_scan_X_14-03-2026.pngS5_scan_Y_14-03-2026.pngCESAR_GUI_10.1.26_14-Mar-26_18-19-38.pngCESAR_GUI_10.1.26_14-Mar-26_18-03-19.pngintensity_scan_pp_14-03-2026.png 
  286   Fri, 13.03.2026 09:43:20 K. Wojcikpp run 2026: preparation   CESAR_GUI_10.1.26_13-Mar-26_09-39-35.pngCESAR_GUI_10.1.26_13-Mar-26_09-52-09.png 
  285   Thu, 12.03.2026 13:26:36 K. Wojcikpp run 2026: CEDAR scanDiaphragm: 0.149 mm

 PRESS       TRIGG          6-FOLD      
 CESAR_GUI_10.1.26_12-Mar-26_13-23-59.png 
  284   Sun, 30.11.2025 12:29:47 K. WojcikPb 2025 run setup[B]Material in counters:[/B]
[LIST]
[*] S1: quartz, 80 x 80 x 2 mm
 10x 
  283   Sat, 29.11.2025 15:29:20 K. WojcikPb 2025 run prepatarionSun, 30.11.2025 12:15:03 We had a lot of tuning
and scans yesterday.
 10x 
  282   Sun, 23.11.2025 13:14:48 K. WojcikS5 beam scanSun, 23.11.2025 13:16:09 S5 X position (field
OFF) -280.9 mm.
 Screenshot_2025-11-23_131527.pngScreenshot_2025-11-23_133121.png 
  281   Sun, 9.11.2025 00:01:17 K. WojcikHL proton run 300 GeV - setup in the middle of the runHV settings:
S1_1, S1_2, S1_3 S1_4: 2200V (negative)
S2: 800 V (negative)
 printscreen_281_20251109-001311.pngprintscreen_281_20251109-001430.pngCESAR_GUI_10.1.11_09-Nov-25_00-35-40.pngCESAR_GUI_10.1.11_08-Nov-25_19-43-57.png 
  280   Sun, 26.10.2025 19:01:28 K. WojcikPreparation for proton run1. S4: I repaired position measure now it
doesn't block and bend anymore. Scintillator
was pushed off the lightguide. I attached
 15x 
  279   Sat, 5.07.2025 12:24:51 K. Wojcikfragmentation run 2025 preparaionsSat,  5.07.2025 12:25:00 I was asked by AD
to lower S4 HV, because maximum threshold
was not high enough. It was 1550, now it
  
  278   Sun, 29.06.2025 13:42:55 Sk SP KWTarget out - in scanPosition  of S2 on beam 85.1 mm

Oxygen run
  
  277   Thu, 19.06.2025 20:31:18 Y. BalkovaHow to calculate beam momentum1. Give Nikos currents on BEND02 and 03, ask
for BL [T*m] ("There is a formula with parameters."
(c) Nikos) Example: for both currents around
  
  276   Tue, 3.06.2025 09:52:39 K. Wojcik, Y. BalkovaBeam tuning @ neutrino run 2025protons 120 GeV/c

S3, S3' and S4 has POSITIVE polarisation
 14x 
  275   Mon, 25.11.2024 11:36:06 Y. Balkova, K. WojcikBeam and counters during fragmentation run 2024Mon, 25.11.2024 setting up for the fragmentation
run
 printscreen_275_20241127-104832.png 
  274   Thu, 21.11.2024 21:48:28 K. WojcikPb run 2024 - Significant beam tuningBeam angle tuned (position on BPD1 moved more
than 10mm).
 V0scan_21nov2024_21-30.pngprintscreen_274_20241121-215152.pngprintscreen_274_20241121-215229.png 
  273   Mon, 4.11.2024 16:49:29 K. WojcikS3 PM replacement and position scanMon,  4.11.2024 16:39:12 After turning on
magnetic field, we lost S3 signal. PM was
replaced for the small one (the same as during
 S3_scan.png 
  272   Wed, 30.10.2024 18:17:03 Y. BalkovaBeam and counters during Pb run 2024Wed, 30.10.2024: Pb run preparation began.

S4 removed from beam.
 printscreen_272_20241102-163218.pngV0_PbPb2024.pngprintscreen_272_20241102-195641.pngV1_PbPb2024.pngprintscreen_272_20241103-011144.png 
  271   Sat, 19.10.2024 19:52:47 Y. BalkovaBeam and counters during pi+C run 202416.10.2024 V1 movement and HV connected and
tested
 21x 
  270   Mon, 22.07.2024 15:02:48 K. WojcikBeam tuning22.07.2024 V0 HV raised to 980V. PP added
active splitter (better signal shape) and
changed discriminator (lower threshold).
  
  270   Mon, 22.07.2024 15:02:48 K. WojcikBeam tuning22.07.2024 V0 HV raised to 980V. PP added
active splitter (better signal shape) and
changed discriminator (lower threshold).
  
  269   Sat, 6.07.2024 14:52:01 Y. BalkovaBeam and counters during neutrino run 2024[B]S1[/B]: Was installed and HV tested. Piotr
and I opened S1 box on evening of July 5th,
because the counts were too low. It appeared
 12x 
  269   Sat, 6.07.2024 14:52:01 Y. BalkovaBeam and counters during neutrino run 2024[B]S1[/B]: Was installed and HV tested. Piotr
and I opened S1 box on evening of July 5th,
because the counts were too low. It appeared
 12x 
  268   Fri, 13.10.2023 22:48:01 K. WojcikBPD2 removedBPD2 removed from the beamline 12-10-2023
at 9:19 am.
  
  267   Sat, 30.09.2023 17:23:19  Pb test settingsChannel 1 = S1
channel 2 = V1
 printscreen_267_20230930-172355.pngprintscreen_267_20231001-122116.pngprintscreen_267_20231003-235407.png 
  266   Tue, 8.08.2023 15:46:09 Y. BalkovaLessons learned[B]Never ever trust anyone again![/B]

When the beam configuration changes:
 8x 
  265   Wed, 2.08.2023 21:16:21 Y.BalkovaBeam tuning p@120 GeVAfter the THC scan, the pressure was set to
0.87 bar.
[LINE]
 7x 
  264   Fri, 28.07.2023 21:26:02 Y. BalkovaFluctuating beam position in YSeveral last days we had a problem with beam
fluctuating in Y direction. First, Nikos
blamed the power supplies for the several
 printscreen_264_20230728-221251.png 
  263   Wed, 26.07.2023 22:41:12 Y. BalkovaBeam tuning K@60 GeVWed, 26.07.2023 22:41:27 YB: After the magnets
were ramped back up and MD was finished,
I began to tune the beam around 19:00. The
 CEDAR_scan_60GeV.pngCEDAR_scan_piK.pngCEDAR_forInaki.png 
  262   Tue, 25.07.2023 18:23:43 K. WojcikCEDAR scan p 120 GeV PRESS       TRIGG          6-FOLD       
  7-FOLD          8-FOLD   
------------------------------------------------------------------------
 cedar120_1.pngcedar120_2.pngcedar_alignment.png 
  261   Mon, 24.07.2023 21:56:24 K. WojcikCedar scan 90GeV/c protons PRESS       TRIGG          6-FOLD       
  7-FOLD          8-FOLD   
------------------------------------------------------------------------
 CEDAR_2023-07-23-protons90GeV.png 
  260   Sun, 23.07.2023 02:37:34 K. WojcikBeam after Nikos tuned it22.07.2023 - Nikos tuned the beam perfectly.
Cedar alignment was done and diaphragm was
set to 3 mm (read value: 2.835). Quick pressure
 CEDAR22_07_23_d2_831.png 
  259   Wed, 19.07.2023 11:15:06 K. WojcikBeam tunningWed, 19.07.2023 11:16:01 After beam retunning
on the accelerator site (wobbling was changed),
we need to retune it as well. Scans started.
 printscreen_259_20230719-204451.pngbeam_counters.jpg 
  258   Tue, 18.07.2023 12:09:40 K. WojcikBeam tuningThe beam has moved this night, so it needs
to be trimmed.
 beamVertical.pngbeamHorizontal.png 
  257   Mon, 17.07.2023 17:15:37 K. WojcikTHC scan - looks betterAfter Inaki's intervention scan looks better.
However, the kaon step is barely visible
Mon, 17.07.2023 20:45:50 I made skanfor Yoshikazu
 thcGood1.pngthcnotsogood.pngthcnotsogood2.pngthc_protons.png 
  256   Fri, 14.07.2023 12:31:35 K. WojcikTHC scan - not successfullFri, 14.07.2023 12:35:11 THC scan finished,
but it is does not look as I expected (thcNew1.png)
<- it was only total efficienty in very wide
 thcNew1.pngthcNew8.pngthcNew8_with_protons.png 
  255   Fri, 7.07.2023 17:26:05 S. PulawskiCedar and THC scans scan 60GeV/cELOG with 2015 trigger monitor screenshots:
[URL=http://na61-elog.cern.ch/Shift+logbook/1423]link[/URL]
 cedar_correct.pngthc1.pngthc2.pngcedarKaonsPions.png 
  254   Fri, 7.07.2023 10:55:59 K. WojcikNeutrino run 2023 - CEDAR scanFri,  7.07.2023 10:56:14 Cedar scan completed
within the range 1.4 - 1.9 bar. Need to extend
the upper limit to 2.1. Starting the new
 CEDAR_scan60gev.png 
  253   Thu, 6.07.2023 09:08:19 K. WojcikNeutrino run 2023 CEDAR scan issueThu,  6.07.2023 09:19:17 Cedar scan not finished
by night. Pressure got stuck at 3.6 bar (scanning
from 3.8 to 1 bar)
 cedar_scan_fail.png 
  252   Wed, 16.11.2022 13:33:24 T.CzopowiczTest entryPiotr asked me to check, weatcher I can ad
stuff here...
  
  251   Fri, 16.11.2018 19:14:19 A. MarcinekPb at 150A GeV/c beam retuning by Nikos : focus in XZAtt. 1 - settings

Att. 2 - 4 - the beam
 CESAR_GUI_8.5.14_16-Nov-18_19-14-46.pngprintscreen_251_20181116-191601.pngprintscreen_251_20181116-191609.pngprintscreen_251_20181116-191616.png 
  250   Tue, 13.11.2018 00:55:50 A. MarcinekPb at 150A GeV/c beam tuningAtt. 1 - 3 - after taking over from Nikos

Att. 4 - 6 - after beam tuning
 7x 
  249   Tue, 6.11.2018 18:30:32 Yuliia Balkova, Kamil WojcikBeam tuning for target-inTue,  6.11.2018 18:32:04 The best beam position
ever in this run
 printscreen_249_20181106-183310.pngprintscreen_249_20181106-183347.png 
  248   Sun, 4.11.2018 12:53:51 K. WojcikRadiation alarmsDue to low beam intensity collimators were
opened last time the beam was tuned. It caused
radiation alarms. Investigating.
 printscreen_248_20181104-135523.pngprintscreen_248_20181104-135539.pngprintscreen_248_20181104-135550.png 
  247   Sun, 4.11.2018 11:48:58 Nikolaos CharitonidisBeam Dears, 

I was called by the CCC that very high radiation
 COLLS.jpgXDWC.jpg 
  246   Sun, 4.11.2018 00:39:35 K. WojcikBeam tiningSun,  4.11.2018 00:47:12 Beam tuned.  printscreen_246_20181104-004745.pngprintscreen_246_20181104-004804.pngprintscreen_246_20181104-004830.png 
  245   Thu, 1.11.2018 19:50:47 Yuliia Balkova, Kamil WojcikBeam tuning 1.11.2018Thu,  1.11.2018 19:51:32 Beam tuning started
Thu,  1.11.2018 21:15:40 I guess Brant is
satisfied
 skan.pngskan1.png 
  244   Mon, 23.07.2018 21:48:01 K.WojcikBeam tunning - noise around beamAfter moving target out there's some "noise"
around beam.
 printscreen_244_20180723-215424.pngprintscreen_244_20180723-215436.pngprintscreen_244_20180723-215449.png 
  243   Thu, 12.07.2018 14:39:51 K.WojcikTunnung beam spot size - successBeam is tuned:
-first absorber is in the beam (XCON.021.026
Pb 4mm)
 printscreen_243_20180712-144555.pngprintscreen_243_20180712-145714.pngprintscreen_243_20180712-145745.pngprintscreen_243_20180712-145807.pngprintscreen_243_20180712-145819.png 
  242   Wed, 11.07.2018 23:52:04 K.WojcikTunnung beam spot sizeWed, 11.07.2018 23:52:10 
I've been doing this for about 3 hours. It
is impossible to make it with collimators
 CESAR_GUI_8.5.7_12-Jul-18_00-02-11.pngprintscreen_242_20180712-000453.pngprintscreen_242_20180712-000459.png 
  241   Sun, 8.07.2018 00:42:15 K. WojcikBad beam after 23:00 7.07.2018Sun,  8.07.2018 00:50:14 
Doors were open. Beam was shifted a little
in YZ plane. Tunning...
 printscreen_241_20180708-024339.pngprintscreen_241_20180708-024408.png 
  240   Sat, 7.07.2018 17:21:54 K. WojcikCEDAR scan h+ beam @120 - USNA61 runPRESS       TRIGG          6-FOLD        
 7-FOLD          8-FOLD   
------------------------------------------------------------------------
 CESAR_GUI_8.5.6_07-Jul-18_17-37-13.png 
  239   Fri, 27.04.2018 11:26:25 E. KapturHadrons at 120 GeV/c - PSD Callibration   printscreen_239_20180427-124554.pngprintscreen_239_20180427-124618.pngprintscreen_239_20180427-124637.pngprintscreen_239_20180427-124709.pngprintscreen_239_20180427-124724.png 
  238   Thu, 26.04.2018 00:26:51 E. KapturHadrons at 80 GeV/c - PSD Callibration   printscreen_238_20180426-002703.pngprintscreen_238_20180426-002719.pngCESAR_GUI_8.4.8_26-Apr-18_00-28-18.png 
  237   Wed, 25.04.2018 23:21:59 E. KapturHadrons at 40 GeV/c - PSD Callibration   printscreen_237_20180425-232218.pngprintscreen_237_20180425-232233.pngCESAR_GUI_8.4.8_25-Apr-18_23-23-35.png 
  236   Thu, 7.12.2017 14:00:57 E. KapturXe+La at 13A GeV/c   printscreen_236_20171207-140128.pngprintscreen_236_20171207-143400.pngCESAR_GUI_8.3.6_07-Dec-17_14-34-51.png 
  235   Wed, 29.11.2017 13:03:55 A. MarcinekXe at 40 AGeV/c : first beam tuningAtt. 1 - 3 - after taking over from Nikos

Att. 4 - Upstream reference beam positions
 9x 
  234   Sat, 25.11.2017 13:07:49 E. Kaptur, A. MarcinekS2 scanS2 to 67.5 mm  S2_scan.png 
  233   Sat, 25.11.2017 11:20:56 E.KapturBeam Spill Structure131
1.0 18.0
2.0 28.0
 spill11.csvspill10.csv 
  232   Tue, 21.11.2017 13:18:11 A. MarcinekXe at 30 AGeV/c : adjustments21.11.2017 about 6:30 AM: After access beam
moved by 8mm on BPD1y. Tuned BEND03 222.6
-> 223.3
  
  231   Mon, 20.11.2017 13:32:03 A. MarcinekXe at 30 AGeV/c : first beam tuningAtt. 1 Upstream reference beam positions

Att. 2 - 4 - After taking over from Nikos
 8x 
  230   Tue, 14.11.2017 00:43:10 A. Marcinek Xe at 19 AGeV/c : day 114.11.2017 00:57:37 AM: BEND03 141.7 -> 141.2
(about 0.4 shift on BPD1)
 6x 
  229   Mon, 13.11.2017 15:51:21 A. MarcinekXe at 19 AGeV/c : first beam tuningAtt 1 - 3: final words of Nikos for now. Maybe
he will still try to move the focus from
BPD2/S2 to the target. Anyway I take over
 11x 
  228   Fri, 10.11.2017 13:13:51 A. MarcinekXe at 75 AGeV/c : day 3 to day 78.11.2017 21:30 AM: BEND03 561.5 -> 561.0,
TRIM05 5 -> 7, BEND6 -483 -> -482
  
  225   Tue, 7.11.2017 02:10:20 A. MarcinekXe at 75 AGeV/c : day 2 7.11.2017 02:10:23 AM: beam moved upstream
causing T4/T1 to raise. BEND03 562.8 -> 562.0
 QAmonitorBPD.pngPrzechwycenie_obrazu_ekranu_2017-11-07_13-52-18.pngPrzechwycenie_obrazu_ekranu_2017-11-07_13-52-31.png 
  224   Mon, 6.11.2017 17:45:40 A. MarcinekXe at 75 AGeV/c : first beam tuningAtt. 1 - beam on beamline chambers (upstream
ones are important as reference)
 6x 
  223   Sun, 5.11.2017 14:09:43 A. MarcinekXe at 150 AGeV/c : day 14 5.11.2017 14:10:49 AM: Realigned on analog
wire chamber upstream: BEND03 1135 -> 1131,
TRIM05 -115 -> -55
  
  222   Tue, 31.10.2017 14:28:06 A. MarcinekXe at 150 AGeV/c : day 91:40 Beam shifted to -0.32, tuned TRIM05 to
-18
  
  221   Mon, 30.10.2017 17:39:44 E. Kaptur, A. MarcinekS2 scanMoved S2 up by 7mm to 62.1mm

[IMG]elog:221/1[/IMG]
 scan.pngprintscreen_221_20171030-190346.png 
  220   Mon, 30.10.2017 15:10:21 A. MarcinekXe at 150 AGeV/c : day 8In a meeting with Karel and Nikos we decided
that the fault is in the beamline
(maybe thermal instability of bends) so we
  
  219   Sun, 29.10.2017 08:41:07 A. MarcinekXe at 150 AGeV/c : day 729.10.2017 08:41:03 AM: The beam is off in
BPD3y: -0.21 and too high in XWCA.021.387
~ 2-5 mm.
 CESAR_GUI_8.3.6_29-Oct-17_15-38-47.png 
  218   Sat, 28.10.2017 14:15:42 A. MarcinekXe at 150 AGeV/c : day 6After many hours managed to retune the beam
and bring back stable conditions.
According to Nikos we should not try to compensate
 9x 
  217   Thu, 26.10.2017 20:32:41 A. MarcinekXe at 150 AGeV/c : day 4Intensity scan att. 1 - we stay at 150k -
below saturation - for safety of VD and to
be sure that we don't screw up PSD measurememt.
 printscreen_217_20171026-203319.pngCESAR_GUI_8.3.6_26-Oct-17_21-08-52.png 
  216   Wed, 25.10.2017 21:48:05 A. MarcinekXe at 150 AGeV/c : day 3Moved beam on the nominal position - att.
1 - 3, (4 & 5 - zooms)
 CESAR_GUI_8.3.6_25-Oct-17_21-51-18.pngprintscreen_216_20171025-215202.pngprintscreen_216_20171025-215209.pngprintscreen_216_20171025-215221.pngprintscreen_216_20171025-215235.png 
  215   Tue, 24.10.2017 21:25:04 A. MarcinekXe at 150 AGeV/c : day 2We realized that the spill structure is ~flat
upstream (XSCI.021.131), but gets worse downstream
(att. 5). So the conclusion was that beam
 6x 
  214   Tue, 24.10.2017 01:26:24 A. MarcinekXe at 150 AGeV/c : initial settingsInitial settings by Nikos - att. 1 - 3 - the
focus is in a wrong place, but the beam is
so narrow that it is irrelevant (for now);
 8x 
  213   Wed, 18.10.2017 18:29:01 A. Marcinekp at 40 GeV/c : new settings from NikosSetting 1 - att. 1 - 4 : disaster

Setting 2 - att. 5 - 8 : another disaster
 16x 
  212   Sat, 14.10.2017 20:54:28 A. Marcinekp at 40 GeV/c : attempt on data taking efficiency optimization x 2 continuedSetting 13 - going back to setting 10 with
opened collimators - att. 1 - 3 - no significant
change in beam shape despite going 10 times
 7x 
  211   Sat, 14.10.2017 17:25:14 A. Marcinekp at 40 GeV/c : attempt on data taking efficiency optimization x 2[COLOR=red]S1 trigger by cable switch + taking
out V1 from CAMAC readout [/COLOR]
 50x 
  210   Fri, 13.10.2017 19:12:00 A. Marcinekp at 40 GeV/c : attempt on data taking efficiency optimizationNew file .201. - att 1, 2

Nikos thought that the beam envelope plots
 9x 
  209   Fri, 13.10.2017 16:52:25 A. Marcinekp at 40 GeV/c : attempt on proton content optimizationAtt. 1 - yesterdays settings, but S1=400k
- i.e. no absorber, default B1, thick primary
target
 16x 
  208   Thu, 12.10.2017 20:48:09 A. Marcinekp at 40 GeV/c: bringing back parallel beamAtt. 1 - last "focused beam" settings save

Moved to "parallel beam" (Att. 2 - 4)
 CESAR_GUI_8.3.6_12-Oct-17_20-45-02.pngCESAR_GUI_8.3.6_12-Oct-17_21-50-27.pngprintscreen_208_20171012-215146.pngprintscreen_208_20171012-215154.png 
  207   Thu, 12.10.2017 15:21:00 A. Marcinekp at 40 GeV/c: trigger ratesAtt. 1 - target IN - T2/T1 = 4.37 %

Att. 2 - target OUT - T2/T1 = 0.92 %
 printscreen_207_20171012-153048.pngprintscreen_207_20171012-154347.png 
  206   Thu, 12.10.2017 09:35:19 A. Marcinekp at 40 GeV/c: CEDAR and trim scansWe have roughly 10% protons, 1.7% kaons and
88% pions.
 Pressure_Scan_History_12-Oct-17_09-34-01.pngCESAR_GUI_8.3.6_12-Oct-17_02-25-35.pngCESAR_GUI_8.3.6_12-Oct-17_14-11-37.png 
  205   Tue, 10.10.2017 21:34:36 A. Marcinekp at 40 GeV/c : attempt 2After long tuning by Nikos to reduce beam
size at our detectors
 printscreen_205_20171010-214827.pngprintscreen_205_20171010-214834.pngprintscreen_205_20171010-214841.pngCESAR_GUI_8.3.6_10-Oct-17_21-51-36.png 
  204   Tue, 10.10.2017 11:13:34 A. Marcinekp at 40 GeV/c : attempt 1First reduced COLLs 7-10 to -5:5 from -10:10
to reduce S1 count from 600k to 160k
 7x 
  203   Fri, 1.09.2017 15:13:37 E. Kapturp at 120 GeV/c   CESAR_GUI_8.2.0_22-Aug-17_15-01-03.pngPressure_Scan_History_22-Aug-17_15-01-13.png 
  202   Fri, 1.09.2017 15:12:03 E. Kapturp at 90 GeV/cpi - 37%   
  201   Wed, 16.08.2017 18:47:36 E. KapturPi- at 60 GeV/cBeam from CERN after angle and position corrections:
Att 1, Att 2.
 10x 
  200   Sat, 5.08.2017 01:54:47 E. Kapturpi+ at 60 GeV/c[IMG]elog:200/2[/IMG][IMG]elog:200/1[/IMG]  CESAR_GUI_8.1.1_05-Aug-17_01-53-31.pngPressure_Scan_History_05-Aug-17_01-53-43.png 
  199   Mon, 19.06.2017 23:11:56 E. Kaptur, S. Pulawskip + Pb @ 30 GeV/c[IMG]elog:199/3[/IMG][IMG]elog:199/2[/IMG][IMG]elog:199/1[/IMG]  CESAR_GUI_8.1.1_19-Jun-17_23-44-15.pngPressure_Scan_History_19-Jun-17_23-44-31.pngCESAR_GUI_8.1.1_25-Jun-17_15-12-48.png 
  198   Tue, 6.12.2016 03:56:03 S.Pulawski, E.KapturPb+Pb at 150A GeV/cIntensity is low.  9x 
ELOG V3.1.3-7933898